What is the x-intercept of y=square root x^2+9?
I squared both sides and ended up with 0=x^2+9
subtracted 9 from both sides .... -9=x^2
Since a negative doesn't go under a radical, do I divide by -1? And if I divide by -1 what do I do with -x^2?

Answer by stanbon(75887)

What is the x-intercept of y=square root x^2+9?
------
Let y = 0
Then sqrt(x^2+9) = 0
Then x^2+9 = 0
x^2 = -9
x = +-3i
-----
There is no x-intercept.
Cheers,
Stan H.
===============